Please help me find the values for the functions below given the following piece wise function.
F(x)=
2x if x <-3
x + 5 if -3 <= x < 5
10-3x if x >= 5


f(8)
f(-8)
f(1)

Answer by stanbon(75887)

1. m=-6 through(-4,1)
You want y=mx+b.
You are given m,x, and y
You need to solve for "b".
1=(-6)(-4)+b
b=-9
Y=(-6)x-9
----------------
2. Runs through (2,3) and(-3,7)
m=[7-3]/[-3-2]=-4/5
You now have m, x, y; so, solve for "b"
3=(-4/5)2+b
3=-8/5 + b
b=23/5
y=(-4/5)x+23/5
Cheers,
Stan H.